# Globalization of Partial Actions of Ordered Groupoids on Rings

### v1

@@ Abstract.
We provide a necessary and sufficient condition to the existence of an ordered globalization of a partial ordered action of an ordered groupoid on a ring and we also present criteria to obtain uniqueness. Furthermore, we apply those results to obtain a Morita context and to show that an inverse semigroup partial action has a globalization (unique up to isomorphism) if, and only if, it is unital.
2020 AMS Subject Classification: Primary. 20L05, 18B40, 20N02. Secondary. 16S35, 20M18
Keywords: ordered groupoid, partial action, ordered globalization, Morita theory, inverse semigroup
@@ 1. Introduction
Much has been developed in the theory of actions and partial actions of groupoids. For instance, Galois theories were considered in [4], [5], [8], [14], [24], [25], [26] and [27], actions in sets and duality theorems in [9], generalizations of classic results of group theory such as isomorphism theorems, Lagrange’s Theorem, Sylow theorems, classification of groupoids and properties of cosets in [1], [2], [6] and [21], and partial actions of ordered groupoids in [3], [15] and [22].
@@
The concept of partial group action was given by Exel in [13] aiming to classify a certain class of $C^{*}$-algebras. J. Kellendonk and M. Lawson [19] investigated applications in several other areas, such as graph immersions and inverse semigroups [19], $\mathbb{R}$-trees [28] and tilings of the Euclidean space [18]. Much more can be found in the surveys written by Dokuchaev in [10] and [11]. The definition of partial action of a groupoid was given by Bagio and Paques in [4]. Among other results, they proved a condition for the existence and uniqueness of a globalization. Partial ordered actions of ordered groupoids on rings were introduced in [3], but no study on the existence of globalization has been done so far.
@@
Our main goal in this work is to analyze the existence and uniqueness of ordered globalizations of partial ordered actions of ordered groupoids on rings. In Section 2 we provide a background about partial ordered actions of ordered groupoids on rings and fix some notations. In Section 3 we prove that a partial ordered action of an ordered groupoid on a ring has a globalization if, and only if, it is unital. In Section 4 we define strong partial ordered actions, pseudoassociative groupoids and minimal globalizations, showing that a unital strong partial ordered action of a pseudoassociative groupoid on a ring always has a unique minimal globalization. Finally, in Section 5 we present two applications: we construct a Morita context between the crossed products of an ordered groupoid acting partially on a ring and its globalization and we apply our results in the case of inverse semigroups acting on rings via the Ehresmann-Schein-Nambooripad (ESN) Theorem [20, Theorem 4.1.8], which states that the category of inverse semigroups and premorphisms is isomorphic to the category of inductive groupoids and ordered functors.
Throughout, rings and algebras are associative and unital unless otherwise stated.
@@ 2. Preliminaries
We recall that a groupoid $\mathcal{G}$ is a small category in which every morphism is an isomorphism. We denote by $\mathcal{G}_{0}$ the set of objects of $\mathcal{G}$. Observe that $\text{id}:\mathcal{G}_{0}\rightarrow\mathcal{G}$, given by $\text{id}(x)=\text{id}_{x}$, is an injective map and whence we identify $\mathcal{G}_{0}\subseteq\mathcal{G}$. Given $g\in\mathcal{G}$, the domain and the range of $g$ will be denoted by $d(g)$ and $r(g)$, respectively. Hence, $d(g)=g^{-1}g$ and $r(g)=gg^{-1}$. For all $g,h\in\mathcal{G}$, we write $\exists gh$ whenever the product $gh$ is defined. We denote $\mathcal{G}_{2}=\{(g,h)\in\mathcal{G}\times\mathcal{G}:\exists gh\}$.
@@
Let $\mathcal{G}$ be an ordered groupoid and $A$ a ring. We say that $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a partial action of $\mathcal{G}$ on $A$ if $A_{r(g)}$ is an ideal of $A$, $A_{g}$ is an ideal of $A_{r(g)}$, $\alpha_{g}:A_{g^{-1}}\to A_{g}$ is an isomorphism of rings, for all $g\in\mathcal{G}$, and
  1. (P1)
    $\alpha_{e}=\text{Id}_{A_{e}}$, for all $e\in\mathcal{G}_{0}$, and $A\subseteq\sum_{e\in\mathcal{G}_{0}}A_{e}$;
  2. (P2)
    $\alpha_{h}^{-1}(A_{g^{-1}}\cap A_{h})\subset A_{(gh)^{-1}}$, for all $(g,h)\in\mathcal{G}_{2}$;
  3. (P3)
@@
The condition $A\subseteq\sum_{e\in\mathcal{G}_{0}}A_{e}$ is required to recover the concept of partial group actions on rings, once if $G$ is a group with identity element $1_{G}$ acting partially on a ring $A$ via $\alpha$, then $A_{1_{G}}=A$ and $\alpha_{1_{G}}=\text{Id}_{A}$. Similar conditions have already appeared in some other papers, such as [9] (on the notion of a “split” $\mathcal{G}$-set) and in (P4) of [23].
By [4, Lemma 1.1], it is also true that:
@@
We say that a partial action $\alpha$ is unital if each $A_{g}$ is a unital ring with central and idempotent identity $1_{g}$. Clearly, $A_{g}=A1_{g}=1_{g}A$, for all $g\in\mathcal{G}$. We say that $\alpha$ is preunital if $A_{e}$ is a unital ring with central and idempotent identity $1_{e}$, for all $e\in\mathcal{G}_{0}$.
Furthermore, $\alpha$ is said to be a partial ordered action, or simply P.O. action, if
  1. (PO)
    If $g\leq h$, then $A_{g}\subseteq A_{h}$ and $\alpha_{g}=\alpha_{h}|_{A_{g^{-1}}}$.
Two P.O. actions $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $A$ and $\gamma=(C_{g},\gamma_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $C$ are isomorphic (we denote by $\alpha\simeq\gamma$) if there is a collection of ring isomorphisms $\{\varphi_{e}:A_{e}\to C_{e}:e\in\mathcal{G}_{0}\}$ such that
@@
linearly extended. We have that $A\ltimes_{\alpha}\mathcal{G}$ is a ring but it is not necessarily associative nor unital. A condition to ensure the associativity of $A\ltimes_{\alpha}\mathcal{G}$ is to ask $\alpha$ to be a unital partial action (c.f. [3, Proposition 3.1] and [12, Corollary 3.2]). Also, by [3, Propostion 3.3], if $\mathcal{G}_{0}$ is finite, then $A\ltimes_{\alpha}\mathcal{G}$ is unital with identity $1_{A\ltimes_{\alpha}\mathcal{G}}=\sum_{e\in\mathcal{G}_{0}}1_{e}$.
@@ 3. Globalization
From now on, assume that all actions and partial actions are preunital, unless stated otherwise.
Let $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be an ordered action of an ordered groupoid $\mathcal{G}$ on a ring $B$. Let $A$ be a unital ideal of $B$. We shall construct a P.O. action $\alpha$ of $\mathcal{G}$ on $A$ by restricting $\beta$. Consider, for each $e\in\mathcal{G}_{0}$, $A_{e}:=A\cap B_{e}$, which is an ideal of $A$. Now, for all $g\in\mathcal{G}$, define $A_{g}:=A_{r(g)}\cap\beta_{g}(A_{d(g)})$, which is an ideal of $A_{r(g)}$ and also define $\alpha_{g}=\beta_{g}|_{A_{g^{-1}}}$, which is a ring isomorphism.
@@
Now we present a new kind of restriction of $\beta$, which generalizes the standard one. Let $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be an ordered global action of $\mathcal{G}$ on a ring $B$. Consider a unital ideal $A_{e}$ of $B_{e}$, for all $e\in\mathcal{G}_{0}$, such that $A_{e}\subseteq A_{f}$ if $e\leq f$ in $\mathcal{G}_{0}$. Then $A_{g}=A_{r(g)}\cap\beta_{g}(A_{d(g)})$ is an ideal of $A_{r(g)}$, for all $g\in\mathcal{G}$. Define $\alpha_{g}=\beta_{g}|_{A_{g^{-1}}}$, for all $g\in\mathcal{G}$. Consider $A$ the (unital) ideal generated by $\bigcup_{e\in\mathcal{G}_{0}}A_{e}$. Since $A_{e}$ is an ideal of $B_{e}$, for all $e\in\mathcal{G}_{0}$, and $\beta$ is unital, then $A_{e}$ is an ideal of $B$, for all $e\in\mathcal{G}_{0}$. Hence $A=\sum_{e\in\mathcal{G}_{0}}A_{e}$ and $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a P.O. action of $\mathcal{G}$ on $A$. We say that $\alpha$ is a restriction of $\beta$ to the ring $A$. Every standard restriction of $\beta$ is a restriction of $\beta$.
Consider now the ideal $C_{g}$ of $B$ generated by $\bigcup_{r(h)\leq r(g)}\beta_{h}(A_{d(h)})$. Since $A_{d(h)}$ is unital and $\beta_{h}$ is an isomorphism, for all $h\in\mathcal{G}$, we have that $C_{g}=\sum_{r(h)\leq r(g)}\beta_{h}(A_{d(h)})$. Notice that $C_{g}\subseteq B_{g}$. Defining $\gamma_{g}=\beta_{g}|_{C_{g^{-1}}}$, we have that $\gamma=(C_{g},\gamma_{g})_{g\in\mathcal{G}}$ is an ordered global action of $\mathcal{G}$ on $C=\sum_{e\in\mathcal{G}_{0}}C_{e}$ such that the restriction of $\gamma$ to the ring $A$ coincides with $\alpha$.
The next example illustrates that there are restrictions which are not standard.
@@
Let $\mathcal{G}=\{m_{0},m_{1},n_{0},n_{1}\}$, where $\mathcal{G}_{0}=\{m_{0},n_{0}\}$, $m_{1}^{2}=m_{0}$, $n_{1}^{2}=n_{0}$, $m_{0}\leq n_{0}$ and $m_{1}\leq n_{1}$. Let $R$ be a commutative ring and $A=Re_{1}\oplus Re_{2}\oplus Re_{3}\oplus Re_{4}$, where the $e_{i}$’s are central idempotents pairwise orthogonal whose sum is $1_{A}$. Define
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a preunital P.O. action of an ordered groupoid $\mathcal{G}$ on a ring $A$. A unital ordered global action $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $B$ is said to be a globalization of $\alpha$ if for all $e\in\mathcal{G}_{0}$ there is an injective homomorphism of rings $\varphi_{e}:A_{e}\to B_{e}$ such that
@@
Now we shall prove that a P.O. action has a globalization if, and only if, it is unital.
@@
=(*$\stackrel{{\scriptstyle(*)}}{{=}}f((g^{-1}g|r(h))h)=f((d(g)|r(h))h)\stackrel{{\scriptstyle(**)}}{{=}}f(r(h)h)=f(h),$
where the equality (***) follows by [20, Proposition 4.1.3] and the equality (***** *) follows since $r(h)\leq d(g)$. Similarly $\gamma_{g}\circ\gamma_{g^{-1}}(f)|_{h}=f(h)$. Thus we have that $\gamma_{g}$ is a ring isomorphism.
@@
for all $f\in F_{h^{-1}}$, $k\in\mathcal{G}_{g}$ and $(g,h)\in\mathcal{G}_{2}$, where the equality (***) follows again by [20, Proposition 4.1.3]. Then $\gamma=(F_{g},\gamma_{g})_{g\in\mathcal{G}}$ is a global action of $\mathcal{G}$ on $\mathcal{F}$.
@@
that is, $\gamma_{h}=\gamma_{g}|_{F_{h^{-1}}}$. The equality (*)(*)( * ) follows from the uniqueness of restrictions, because $(h^{-1}|r(k))\leq h^{-1}\leq g^{-1}$ and $d(h^{-1}|r(k))=r(k)$. Hence $\gamma$ is an ordered global action of $\mathcal{G}$ on $\mathcal{F}$.
@@
Let $B_{g}$ be the ideal generated by $\bigcup_{r(h)\leq r(g)}\gamma_{h}(\varphi_{d(h)}(A_{d(h)}))$, for all $g\in\mathcal{G}$. Notice that $B_{g}\subseteq F_{g}$. Consider $B=\sum_{e\in\mathcal{G}_{0}}B_{e}=\left\langle\bigcup_{e\in\mathcal{G}_{0}}B_{e}\right\rangle$. Define $\beta_{g}:=\gamma_{g}|_{B_{g^{-1}}}$, for all $g\in\mathcal{G}$. By construction, $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ is an ordered global action of $\mathcal{G}$ on $B$. The next step is to show that $\beta$ is a globalization of $\alpha$.
@@
=(*)$\stackrel{{\scriptstyle(*)}}{{=}}\begin{cases}\alpha_{k^{-1}h}(a1_{h^{-1}k})\alpha_{k^{-1}h}(1_{h^{-1}k}1_{h^{-1}})\alpha_{k^{-1}}(b1_{k}),\text{ if }r(k)=r(h)=e,\\ 0,\text{ otherwise}\end{cases}$
@@
where (*)(*)( * ) holds because Proposition (i)(ii) implies that $\alpha_{k^{-1}h}(A_{h^{-1}k}\cap A_{h^{-1}})=A_{k^{-1}h}\cap A_{k^{-1}}$.

### v2

@@ Abstract.
We provide a necessary and sufficient condition to the existence of an ordered globalization of a partial ordered action of an ordered groupoid on a ring, and we also present criteria to obtain uniqueness. Furthermore, we apply those results to obtain a Morita context and to show that an inverse semigroup partial action has a globalization (unique up to equivalences) if and only if it is unital.
Key words and phrases:
ordered groupoid, partial action, ordered globalization, Morita theory, inverse semigroup.
2020 Mathematics Subject Classification:
Primary. 20L05. Secondary. 18B40, 20N02, 16S35, 20M18
Corresponding author
@@ 1. Introduction
Much has been developed in the theory of actions and partial actions of groupoids. For instance, Galois theories were considered in [5], [6], [10], [18], [31], [32], [33] and [34], actions in sets and duality theorems in [11], generalizations of classic results of group theory such as isomorphism theorems, Lagrange’s Theorem, Sylow theorems, classification of groupoids and properties of cosets in [2], [3], [8] and [28], and partial actions of ordered groupoids and categories in [4], [20] and [29].
@@
The concept of partial group action was introduced by R. Exel in [15] in order to classify a certain class of $C^{*}$-algebras. J. Kellendonk and M. Lawson [24] investigated applications in several other areas, such as graph immersions and inverse semigroups [24], $\mathbb{R}$-trees [35] and tilings of the Euclidean space [23]. Much more about this theory and applications can be read in Exel’s book [16], and also in the surveys written by M. Dokuchaev in [12] and [13], and by E. Batista in [7], in the later especially in relation to Hopf algebras and weak Hopf algebras.
The definition of partial action of a groupoid was given by D. Bagio and A. Paques in [5]. Among other results, they proved a condition for the existence and uniqueness of a globalization. Partial ordered actions of ordered groupoids on rings were introduced in [4], but so far no study on the existence of globalization has been done.
The globalization problem has been investigated in several other contexts. For instance, in [14] it was proved that a partial action of a group on a unital algebra is globalizable if and only if each ideal of the action is unital, and it is unique up to equivalence. For categories acting partially on sets, a globalization is not unique, but a globalization with an additional universal property, called universal globalization, is unique up to equivalence, and such a globalization always exists [29, Theorem 4]. For semigroups acting partially on sets, there exist two non-isomorphic universal globalizations [25]. For Hopf algebras acting partially on a unital algebra, a globalization is not unique, but a globalization with an additional minimal property, called minimal globalization, is [1, Theorem 3].
@@
The problem of determining when a partial action can be realized as a restriction of a global one is very relevant, because it allows us to undestand how the partial theory behaves compared to the global one. The issue of uniqueness is particularly crucial, as it establishes a well-defined framework for transitioning to global actions. Thus, the primary aim of this work is to analyze the existence and uniqueness of ordered globalizations of partial ordered actions of ordered groupoids on rings. In Section 2, we provide a background on the partial ordered actions of ordered groupoids and establish some notation. In Section 3, we prove that a partial ordered action of an ordered groupoid on a ring has a globalization if and only if it is unital. Section 4 introduces strong partial ordered actions, pseudoassociative groupoids and minimal globalizations, showing that a unital strong partial ordered action of a pseudoassociative groupoid on a ring always has a unique minimal globalization. Finally, in Section 5, we present two applications: we construct a Morita context between the crossed products of an ordered groupoid acting partially on a ring and its globalization, and we apply our results to the case of inverse semigroups acting on rings via the Ehresmann-Schein-Nambooripad (ESN) Theorem.
Throughout, rings and algebras are associative unless otherwise stated.
@@ 2. Preliminaries
We recall that a groupoid $\mathcal{G}$ is a small category in which every morphism is an isomorphism. We denote by $\mathcal{G}_{0}$ the set of objects of $\mathcal{G}$. Observe that $\text{id}:\mathcal{G}_{0}\rightarrow\mathcal{G}$, given by $\text{id}(x)=\text{id}_{x}$, is an injective map, from which we identify $\mathcal{G}_{0}\subseteq\mathcal{G}$. Given $g\in\mathcal{G}$, the domain and the range of $g$ will be denoted by $d(g)$ and $r(g)$, respectively. Hence, $d(g)=g^{-1}g$ and $r(g)=gg^{-1}$. For all $g,h\in\mathcal{G}$, we write $\exists gh$ whenever the product $gh$ is defined. We denote $\mathcal{G}_{2}=\{(g,h)\in\mathcal{G}\times\mathcal{G}:\exists gh\}$.
@@
Let $\mathcal{G}$ be a groupoid and $A$ be a ring. We define $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ as a partial groupoid action of $\mathcal{G}$ on $A$ if:
1.$A_{r(g)}\triangleleft A$, $A_{g}\triangleleft A_{r(g)}$, and $\alpha_{g}:A_{g^{-1}}\to A_{g}$ is an isomorphism of rings, for all $g\in\mathcal{G}$;
2. The following conditions hold:
  • (P1)
    $\alpha_{e}=\text{Id}_{A_{e}}$, for all $e\in\mathcal{G}_{0}$, and $A=\sum_{e\in\mathcal{G}_{0}}A_{e}$;
  • (P2)
    $\alpha_{h}^{-1}(A_{g^{-1}}\cap A_{h})\subseteq A_{(gh)^{-1}}$, for all $(g,h)\in\mathcal{G}_{2}$;
  • (P3)
@@ 2. The following conditions hold:
Furthermore, a partial action $\alpha$ is said to be a partial ordered action, or simply P.O. action, if the following condition holds:
  1. (PO)
    If $g\leq h$, then $A_{g}\subseteq A_{h}$ and $\alpha_{g}=\alpha_{h}|_{A_{g^{-1}}}$.
Remark 2.2.
We highlight some facts concerning the definition of partial action of a groupoid.
  • 1.
    The concept of a partial groupoid action was first defined by Bagio and Paques in [5], although the authors did not require the condition $A=\sum_{e\in\mathcal{G}_{0}}A_{e}$. We incorporate it into the definition in order to recover the concept of partial group actions on rings.
  • 2.
    A similar, but stronger, condition has already appeared in [30]. The property (P4) of [30] requires $A=\displaystyle\bigoplus_{e\in\mathcal{G}_{0}}A_{e}$. This type of action was later called orthogonal partial action [27, Section 2]. In a P.O. action of an ordered groupoid, the orthogonality is translated into $A=\displaystyle\bigoplus_{e\in\text{max }\mathcal{G}_{0}}A_{e}$, where max $\text{max }\mathcal{G}_{0}$ is the set of maximal elements in $\mathcal{G}_{0}$ relative to the order of $G$. This means that given $f_{1},f_{2}\in\text{max }\mathcal{G}_{0}$ such that $A_{f_{1}}\neq A_{f_{2}}$, for each $e\in\mathcal{G}_{0}$ such that $e\leq f_{1},f_{2}$, we have $A_{e}=0$. Therefore, this condition is quite restrictive in the ordered case, and that is why we are asking for a weaker condition.
By [5, Lemma 1.1], it is also true that:
@@
A unital partial action of $\mathcal{G}$ on $A$ is a partial action $\alpha$ such that every $A_{g}$, for $g\in\mathcal{G}$, is generated by a central idempotent $1_{g}$ of $A$. A ring $A$ has local units if there is a set $E\subseteq A$ of central idempotents such that for every $a\in A$ there is $b\in E$ such that $ba=a$. We say that $\alpha$ is preunital if $A_{e}$ is a unital ring with central and idempotent identity $1_{e}$, for all $e\in\mathcal{G}_{0}$. Notice that the condition (P1) implies that if there is a preunital partial action of $\mathcal{G}$ on $A$, then $A$ has local units, but it is not necessarily unital unless $\mathcal{G}_{0}$ is a finite set.
Two P.O. actions $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $A$ and $\gamma=(C_{g},\gamma_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $C$ are equivalent (we denote by $\alpha\simeq\gamma$) if there is a collection of ring isomorphisms $\{\varphi_{e}:A_{e}\to C_{e}:e\in\mathcal{G}_{0}\}$ such that
@@
linearly extended. We have that $A\ltimes_{\alpha}\mathcal{G}$ is a ring, but it is not necessarily associative nor unital. A condition to ensure the associativity of $A\ltimes_{\alpha}\mathcal{G}$ is to ask $\alpha$ to be a unital partial action (c.f. [4, Proposition 3.1] and [14, Corollary 3.2]). Also, by [4, Propostion 3.3], if $\mathcal{G}_{0}$ is finite and $\alpha$ is preunital, then $A\ltimes_{\alpha}\mathcal{G}$ is unital with identity $1_{A\ltimes_{\alpha}\mathcal{G}}=\sum_{e\in\mathcal{G}_{0}}1_{e}\delta_{e}$. If $\alpha$ is preunital and $\mathcal{G}_{0}$ is not finite, then $A\ltimes_{\alpha}\mathcal{G}$ is not unital in general, but it has local units, since $A$ has.
@@ 3. Globalization
Let $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be an ordered action of an ordered groupoid $\mathcal{G}$ on a ring $B$. Let $A$ be an ideal of $B$. We shall construct a P.O. action $\alpha$ of $\mathcal{G}$ on $A$ by restricting $\beta$. Consider, for each $e\in\mathcal{G}_{0}$, $A_{e}:=A\cap B_{e}$, which is an ideal of $A$ and $B$. Now, for all $g\in\mathcal{G}$, define $A_{g}:=A_{r(g)}\cap\beta_{g}(A_{d(g)})$, which is an ideal of $A_{r(g)}$, and also define $\alpha_{g}=\beta_{g}|_{A_{g^{-1}}}$, which is a ring isomorphism.
@@
This fact inspires our next definition.
Definition 3.3.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be an P.O. action of $\mathcal{G}$ on a ring $A$. We say that $\alpha$ is a strong P.O. action if $A_{(e|g)}=A_{e}\cap A_{g}$, for all $g\in\mathcal{G}$, $e\leq r(g)$.
Now we present a new kind of restriction of $\beta$, which generalizes the standard one. Let $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be an ordered global action of $\mathcal{G}$ on a ring $B$. Consider $A_{e}\subseteq B_{e}$ an ideal of $B$, for all $e\in\mathcal{G}_{0}$, such that $A_{e}\subseteq A_{f}$ if $e\leq f$ in $\mathcal{G}_{0}$. Then $A_{g}=A_{r(g)}\cap\beta_{g}(A_{d(g)})$ is an ideal of $A_{r(g)}$, for all $g\in\mathcal{G}$. Define $\alpha_{g}=\beta_{g}|_{A_{g^{-1}}}$, for all $g\in\mathcal{G}$. Consider $A=\sum_{e\in\mathcal{G}_{0}}A_{e}$, which is an ideal of $B$. Hence $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a P.O. action of $\mathcal{G}$ on $A$. We say that $\alpha$ is a restriction of $\beta$ to the ring $A$.
Every standard restriction of an ordered global action is a restriction. The next example illustrates that there are restrictions that are not standard.
@@
Let $\mathcal{G}=\{m_{0},m_{1},n_{0},n_{1}\}$, where $\mathcal{G}_{0}=\{m_{0},n_{0}\}$, $m_{1}^{2}=m_{0}$, $n_{1}^{2}=n_{0}$, $m_{0}\leq n_{0}$ and $m_{1}\leq n_{1}$. Let $R$ be a commutative ring and $A=Re_{1}\oplus Re_{2}\oplus Re_{3}\oplus Re_{4}$, where the $e_{i}$’ s are pairwise orthogonal central idempotents whose sum is $1_{A}$. Define
@@
It is always possible to obtain many P.O. actions from ordered global actions using restrictions. The question we want to answer is that when we can, starting from a P.O. action $\alpha$, give rise to a global ordered action $\beta$ such that a restriction of $\beta$ is equivalent to $\alpha$. We call such an ordered global action of ordered globalization.
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a P.O. action of an ordered groupoid $\mathcal{G}$ on a ring $A$. An ordered global action $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $B$ is said to be a globalization of $\alpha$ if for all $e\in\mathcal{G}_{0}$ there is an injective homomorphism of rings $\varphi_{e}:A_{e}\to B_{e}$ such that
@@
Next, we shall prove that a preunital P.O. action has an ordered globalization if and only if it is unital.
@@
=$\stackrel{{\scriptstyle(*)}}{{=}}f((g^{-1}g|r(h))h)=f((d(g)|r(h))h)\stackrel{{\scriptstyle(**)}}{{=}}f(r(h)h)=f(h),$
where the equality (*) follows by [26, Proposition 4.1.3] and the equality (**∗ ∗) follows since $r(h)\leq d(g)$. Similarly $\gamma_{g}\circ\gamma_{g^{-1}}(f)|_{h}=f(h)$. Thus we have that $\gamma_{g}$ is a ring isomorphism.
@@
for all $f\in F_{h^{-1}}$, $k\in\mathcal{G}_{g}$ and $(g,h)\in\mathcal{G}_{2}$, where the equality (*) follows again by [26, Proposition 4.1.3]. Then $\gamma=(F_{g},\gamma_{g})_{g\in\mathcal{G}}$ is a global action of $\mathcal{G}$ on $\mathcal{F}$.
@@
that is, $\gamma_{h}=\gamma_{g}|_{F_{h^{-1}}}$. The equality ()(*)( ∗ ) follows from the uniqueness of restrictions, because $(h^{-1}|r(k))\leq h^{-1}\leq g^{-1}$ and $d(h^{-1}|r(k))=r(k)$. Hence $\gamma$ is an ordered global action of $\mathcal{G}$ on $\mathcal{F}$.
@@
Let $B_{g}$ be the subring generated by $\bigcup_{r(h)\leq r(g)}\gamma_{h}(\varphi_{d(h)}(A_{d(h)}))$, for all $g\in\mathcal{G}$. Notice that $B_{g}\subseteq F_{g}$. Consider $B=\sum_{e\in\mathcal{G}_{0}}B_{e}$. Define $\beta_{g}:=\gamma_{g}|_{B_{g^{-1}}}$, for all $g\in\mathcal{G}$. Then $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ is an ordered global action of $\mathcal{G}$ on $B$. Observe that $\beta$ is not a unital action in general. The next step is to show that $\beta$ is a globalization of $\alpha$.
@@
=()$\stackrel{{\scriptstyle(*)}}{{=}}\begin{cases}\alpha_{k^{-1}h}(a1_{h^{-1}k})\alpha_{k^{-1}h}(1_{h^{-1}k}1_{h^{-1}})\alpha_{k^{-1}}(b1_{k}),\text{ if }r(k)=r(h)=e,\\ 0,\text{ otherwise}\end{cases}$
@@
where ()(*)( ∗ ) holds because Proposition (i)(ii) implies that $\alpha_{k^{-1}h}(A_{h^{-1}k}\cap A_{h^{-1}})=A_{k^{-1}h}\cap A_{k^{-1}}$.
@@
Remark 3.7.
  • (i)
    The ring $B$ constructed above does not need to have an identity element. Ferrero and Lazzarin have determined in [17] necessary and sufficient conditions on the partial group action $\alpha$ for the ring $B$ to be unital. In this scenario, they called $\alpha$ as finite type.
  • (ii)
    Even though the ring $B$ is not unital, it has local units by construction, and hence it is an idempotent ring, that is, $B^{2}=B$.